Solve the problem. The perimeter of a rectangle is 62 m. If the width were doubled and the length were increased by 21 m, the perimeter would be 120 m. What are the length and width of the rectangle?
Answer by edjones(8007)

A) 2L+2w=62
B) 2(L+21)+2*2w=120
.
B)
2L+42+4w=120
2L+4w+42-42=120-42
.
2L+4w=78
2L+2w=62
---------Subtract
===>2w=16
2w/2=16/2
w=8 m
.
A)
2L+16=62
2L+16-16=62-16
2L=46
2L/2=46/2
L=23 m
.
Ed